# Conditional Generative Modeling for Enhanced Credit Risk Management in Supply Chain Finance

### v1

@@ 3.3.1 Asymptotic Analysis for $r_{1}(l)$ Estimation
    For any given interval $[\tau_{\ell},\tau_{u}]\subset(0,1)$ with $10<\tau_{\ell}<\tau_{u}<$ 1, the density functions $f_{Y}(\cdot|\mathbf{x})$ are uniformly bounded away from 00 and $\infty$ in the sense that
@@
    For any given interval $[\tau_{\ell},\tau_{u}]\subset(0,1)$ with $10<\tau_{\ell}<\tau_{u}<$ 1, there exist positive definite matrices $D_{0}$ and $D_{1}(\tau)$ such that
@@
Assumption 2 and Assumption 3 are commonly adopted in quantile regression analyses to ensure robust and reliable statistical inferences. These conditions facilitate the estimation of quantile regression parameters and support the theoretical underpinnings of asymptotic analyses.
With the above assumptions, the following uniform convergence of $F_{\hat{Y}}(y|\mathbf{x})$ can be obtained.
Lemma 1.
@@
    For any $\mathbf{x}\in\mathcal{X}$, $\hat{Y}(\mathbf{x})\Rightarrow Y(\mathbf{x})$ as $n,m\to\infty$.
@@
where $\Rightarrow$” denotes convergence in distribution.
Asymptotics for $r_{1}(l)$ Estimation
Recall that $\hat{r}_{1}(l)$ and $r_{1}(l)$ represent the estimated and true probabilities of default, respectively. Their difference is precisely:
$\bigl{|}\hat{r}_{1}(l)-r_{1}(l)\bigr{|}\;=\;\Bigl{|}\,F_{\hat{Y}}\bigl{(}\,l/r\!\mid\!\mathbf{x}\bigr{)}\;-\;F_{Y}\bigl{(}\,l/r\!\mid\!\mathbf{x}\bigr{)}\Bigr{|}.$
Hence, by Lemma 1, we obtain:
Theorem 1.
Under Assumptions 13, the following results hold:
  1. (1.a)
    For any $\mathbf{x}\in\mathcal{X}$, and for any $l\in[0,\bar{l}]$, we have $\hat{r}_{1}(l)\to r_{1}(l)$ as $n,m\to\infty$.
  2. (1.b)
    For any $\mathbf{x}\in\mathcal{X}$, we have $\sup\limits_{l\in[0,\bar{l}]}\left|\hat{r}_{1}(l)-r_{1}(l)\right|\to 0$ as $n,m\to\infty$.
These results ensure $\hat{r}_{1}(l)$’s uniform convergence across $l\in[0,\bar{l}]$, supporting adaptive loan sizing in SCF. This analysis not only offers a rigorous foundation for practitioners to trust QRGMM-generated PD estimation but also lays the groundwork for analyzing more complex risk measures like $r_{2}(l)$ and $r_{3}(l)$, explored in subsequent subsections.
3.3.2 Asymptotic Analysis for $r_{2}(l)$ Estimation
In this subsection, we delve deeper into the analysis of $r_{2}(l)=\mathbb{E}\left[(l-rY)^{+}\right],$ and its approximation via QRGMM. Compared to the estimation of the default probability $r_{1}(l)$, which involves the direct evaluation of the estimated conditional distribution function $F_{\hat{Y}}(y|\mathbf{x})$, the estimation of $r_{2}(l)$ is more subtle. It requires analyzing the convergence of the expectation of a transformation of the random variable $Y(\mathbf{x})$, namely the positive part $(l-rY)^{+}$. This complicates the asymptotic analysis since standard weak convergence results do not automatically guarantee convergence of expectations for general functions.
Fortunately, the structure of $r_{2}(l)$ offers an advantage: for all $l\in[0,\bar{l}]$, the function $y\mapsto(l-ry)^{+}$ is uniformly bounded by $\bar{l}$ and continuous in $y$. These properties enable us to invoke a classical result from probability theory regarding the convergence of expectations under weak convergence:
Lemma 2 (Portmanteau Theorem).
Let $Z_{n}\Rightarrow Z$ as $n\to\infty$. Then for any bounded and continuous function $h$, it holds that $\mathbb{E}\left[h(Z_{n})\right]\to\mathbb{E}\left[h(Z)\right].$
This Portmanteau theorem (see, e.g., Van der Vaart (2000, page 6)) provides a critical tool in establishing uniform convergence of $\hat{r}_{2}(l)$. Specifically, for each $l\in[0,\bar{l}]$, we define the transformation $f_{l}(y):=(l-ry)^{+},$ which is uniformly bounded above by $\bar{l}$ and continuous on $[0,\infty)$. Given that $\hat{Y}(\mathbf{x})\Rightarrow Y(\mathbf{x})$ as $n,m\to\infty$ by Lemma (1.a), applying Lemma 2 ensures pointwise convergence $\hat{r}_{2}(l)\to r_{2}(l)$ for each $l$. Furthermore, leveraging the equicontinuity of $f_{l}(y)=(l-ry)^{+}$ with respect to $l$, we can extend this result to uniform convergence over $l\in[0,\bar{l}]$. The result is summarized in the following theorem.
Theorem 2.
Suppose that Assumptions 13 hold, if $m$ satisfies $\frac{1}{m}\to 0$ and $\frac{n}{m}\to\infty$ as $n\to\infty$, then, for any given $\mathbf{x}$, as $m,n\to\infty$, we have
$\sup\limits_{l\in[0,\bar{l}]}\bigl{|}\hat{r}_{2}(l)-r_{2}(l)\bigr{|}\to 0.$
Proof.
For any $l\in[0,\bar{l}]$, the function $f_{l}(y)=(l-ry)^{+}$ is Lipschitz continous in $y$ with Lipschitz constant $r$, and uniformly bounded with $0\leq f_{l}(y)\leq\bar{l}$ for all $y\in\mathbb{R}$. Additionally, for all $y\in\mathbb{R}$ and $l,l^{\prime}\in[0,\bar{l}]$, we have
$|f_{l}(y)-f_{l^{\prime}}(y)|\leq|l-l^{\prime}|,$
which implies that $\{f_{l}\}$ is equicontinuous in the parameter $l$ (since each $f_{l}(y)$ is Lipschitz continuous in $l$ with a common Lipschitz constant constant $1$ regardless of $y$).
For any $\mathbf{x}\in\mathcal{X}$, since $\hat{Y}(\mathbf{x})\Rightarrow Y(\mathbf{x})$ as $n,m\to\infty$ by Lemma (1.a) and each $f_{l}$ is continuous and bounded, by Lemma 2, it follows that
$\hat{r}_{2}(l)=\mathbb{E}[f_{l}(\hat{Y}(\mathbf{x}))]\to r_{2}(l)=\mathbb{E}[f_{l}(Y(\mathbf{x}))]$
for every fixed $l\in[0,\bar{l}]$. To strengthen this to uniform convergence over $l\in[0,\bar{l}]$, fix any $\epsilon>0$. By the compactness of $[0,\bar{l}]$ and the equicontinuity of $\{f_{l}\}$ in $l$, there exists a finite $\delta$-net $\{l_{1},\dots,l_{T}\}\subset[0,\bar{l}]$ such that for any $l\in[0,\bar{l}]$, there exists $j\in\{1,\dots,T\}$ with $|l-l_{j}|<\delta$, where $\delta:=\epsilon/3$.
Then for such $l$ and corresponding $l_{j}$, we decompose the total error as:
$\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l}(Y)]\right|$
$\leq$ $\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\right|$
$+$ $\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\right|$
$+$ $\left|\mathbb{E}[f_{l_{j}}(Y)]-\mathbb{E}[f_{l}(Y)]\right|.$
The first and third terms are bounded by
$\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\right|$
$=$ $\left|\mathbb{E}[f_{l}(\hat{Y}_{n})-f_{l_{j}}(\hat{Y}_{n})]\right|$
$\leq$ $\left|\mathbb{E}[l-l_{j}]\right|=|l-l_{j}|<\delta=\epsilon/3,$
and similarly for $\left|\mathbb{E}[f_{l_{j}}(Y)]-\mathbb{E}[f_{l}(Y)]\right|$.
For the second term, since the set $\{l_{1},\dots,l_{T}\}$ is finite and $\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]\to\mathbb{E}[f_{l_{j}}(Y)]$ for each $j$, there exists $M$ and $N$ such that for all $m>M$, $n>N$ and all $j$,
$\left|\mathbb{E}[f_{l_{j}}(\hat{Y}_{n})]-\mathbb{E}[f_{l_{j}}(Y)]\right|<\epsilon/3.$
Combining the above, we conclude that for all $l\in[0,\bar{l}]$ and all $m,n$ sufficiently large,
$\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l}(Y)]\right|<\epsilon.$
Therefore,
$\sup\limits_{l\in[0,\bar{l}]}\bigl{|}\hat{r}_{2}(l)-r_{2}(l)\bigr{|}=\sup_{l\in[0,\bar{l}]}\left|\mathbb{E}[f_{l}(\hat{Y}_{n})]-\mathbb{E}[f_{l}(Y)]\right|\to 0.$
This asymptotic guarantee is particularly significant in the SCF setting, where the continous and bounded structure of $f_{l}(y)$ simplify the analysis. Without these properties, expectation and tail behavior could complicate the convergence analysis, requiring stronger assumptions. Our approach effectively exploits this structure, leading to the uniform convergence for $r_{2}(l)$ estimation. This result provides theoretical validation for the use of QRGMM in estimating expected loss risk measures, reinforcing its applicability in SCF risk management.
3.3.3 Asymptotic Analysis for $r_{3}(l)$ Estimation
Here, ignoring the MC error, we prove the uniform convergence of the estimated risk measure $\hat{r}_{3}(l)=\mathbb{E}[g_{l}(\hat{Y}(\mathbf{x}))\cdot\mathbf{I}_{\{\hat{Y}(\mathbf{x})<a(l)\}}]$, computed using the QRGMM-based generative model $\hat{Y}(\mathbf{x})$, to the true $r_{3}(l)$. Unlike $r_{2}(l)$, where the function $(l-ry)^{+}$ is continuous and bounded, enabling the use of the Portmanteau Theorem to establish convergence, the general form of $g_{l}(y)\cdot\mathbf{I}_{\{y<a(l)\}}$ poses a challenge. The indicator function may introduce the discontinuity, rendering the transformation non-continuous. Thus, theoretical analysis method applied to $\hat{r}_{2}(l)$ is not directly applicable to $\hat{r}_{3}(l)$
To address this, we take a different approach to establish the uniform convergence for $\hat{r}_{3}(l)$, which is consequently also applicable to $\hat{r}_{1}(l)$ and $\hat{r}_{2}(l)$. Specifically, we first impose the following regularity conditions on $g_{l}(\cdot)$ and $a(l)$:
@@ 3.3.3 Asymptotic Analysis for $r_{3}(l)$ Estimation
    $a(l)$ is monotone on $[0,\bar{l}]$, satisfies $0\leq a(l)\leq a(\bar{l})<\infty$, with $a(l)\to 0$ as $l\to 0$.
@@
    $g_{l}(\cdot)$ is differentiable and Lipschitz continuous with constant $K$ within $y\in[0,a(\bar{l})]$ uniformly over $l\in[0,\bar{l}]$, and uniformly bounded by $g_{\bar{l}}(0)<\infty$.
Assumption (4.a) requires $a(l)$, a threshold function, to be monotonic in $l$ and and vanish as $l\to 0$, reflecting the intuition that smaller loans imply lower trigger levels. Assumption (4.b) ensures that $g_{l}(\cdot)$, representing loss in SCF, is differentiable, Lipschitz continuous with constant $K$ (implying $|g_{l}^{\prime}(y)|\leq K$) within $y\in[0,a(\bar{l})]$, and bounded by $M=g_{\bar{l}}(0)$, reflecting the maximum loss in the worst case that sales $Y=0$ and loan $l=\bar{l}$. These conditions, intuitive and reasonable for SCF risk management, ensure tractable analysis of $\hat{r}_{3}(l)$, leading to the following result.
@@
Under Assumptions 14, if $m$ satisfies $\frac{1}{m}\to 0$ and $\frac{n}{m}\to\infty$ as $n\to\infty$, then for any given $\mathbf{x}$, as $n,m\to\infty$, we have
@@
Recall that
$r_{3}(l)=\mathbb{E}\left[g_{l}(Y(\mathbf{x}))\cdot\mathbf{I}_{\{Y(\mathbf{x})<a(l)\}}\right],$
$\hat{r}_{3}(l)=\mathbb{E}\left[g_{l}(\hat{Y}(\mathbf{x}))\cdot\mathbf{I}_{\{\hat{Y}(\mathbf{x})<a(l)\}}\right].$
Express the expectations:
[…]
The difference is:
$|\hat{r}_{3}(l)-r_{3}(l)|=\left|\int_{0}^{a(l)}g_{l}(y)\,d(F_{\hat{Y}}-F_{Y})(y|\mathbf{x})\right|.$
As $g_{l}(y)$ is differentiable and Lipschitz continuous with constant $K$ within $y\in[0,a(\bar{l})]$ uniformly over $l\in[0,\bar{l}]$, we have $|g_{l}^{\prime}(y)|\leq K$ within $y\in[0,a(\bar{l})]$. Let
$\delta_{n,m}=\sup_{y\geq 0}|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|,$
and denote $M=g_{\bar{l}}(0)$.
For any $\mathbf{x}\in\mathcal{X}$ and $l\in[0,\bar{l}]$, compute the difference:
$|\hat{r}_{3}(l)-r_{3}(l)|$
$=$ $\left|\int_{0}^{a(l)}g_{l}(y)\,dF_{\hat{Y}}(y|\mathbf{x})-\int_{0}^{a(l)}g_{l}(y)\,dF_{Y}(y|\mathbf{x})\right|$
$=$ $\left|\int_{0}^{a(l)}g_{l}(y)\,d(F_{\hat{Y}}-F_{Y})(y|\mathbf{x})\right|.$
Using integration by parts:
@@
$=$ $\left[g_{l}(y)\left(F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})\right)\right]\big{|}_{0}^{a(l)}$
[…]
Evaluate the boundary term:
$\left[g_{l}(y)\left(F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})\right)\right]\big{|}_{0}^{a(l)}$
$=$ $g_{l}(a(l))\left(F_{\hat{Y}}(a(l)|\mathbf{x})-F_{Y}(a(l)|\mathbf{x})\right)$
-- $g_{l}(0)\left(F_{\hat{Y}}(0|\mathbf{x})-F_{Y}(0|\mathbf{x})\right).$
Thus:
$\hat{r}_{3}(l)-r_{3}(l)$
$=$ $g_{l}(a(l))\left(F_{\hat{Y}}(a(l)|\mathbf{x})-F_{Y}(a(l)|\mathbf{x})\right)$
-- $g_{l}(0)\left(F_{\hat{Y}}(0|\mathbf{x})-F_{Y}(0|\mathbf{x})\right)$
[…]
For the boundary terms, since $|g_{l}(a(l))|\leq M$, $|g_{l}(0)|\leq M$, and $|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|\leq\delta_{n,m}$, we have
$|g_{l}(a(l))|\cdot|F_{\hat{Y}}(a(l)|\mathbf{x})-F_{Y}(a(l)|\mathbf{x})|\leq M\delta_{n,m},$
$|g_{l}(0)|\cdot|F_{\hat{Y}}(0|\mathbf{x})-F_{Y}(0|\mathbf{x})|\leq M\delta_{n,m}.$
For the integral term, since $|g_{l}^{\prime}(y)|\leq K$ and $|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|\leq\delta_{n,m}$, we have
$\int_{0}^{a(l)}|F_{\hat{Y}}(y|\mathbf{x})-F_{Y}(y|\mathbf{x})|\cdot|g_{l}^{\prime}(y)|\,dy$
[…]
as $a(l)\leq a(\bar{l})$. Combine them, we have
$|\hat{r}_{3}(l)-r_{3}(l)|$
$\leq$ $M\delta_{n,m}+M\delta_{n,m}+K\delta_{n,m}a(\bar{l})$
@@
Since the bound is independent of $l$, it holds uniformly:
$\sup_{l\in[0,\bar{l}]}|\hat{r}_{3}(l)-r_{3}(l)|\leq(2M+Ka(\bar{l}))\delta_{n,m}.$
By Lemma 1(b), we have $\delta_{n,m}\to 0$, as $n,m\to\infty$. Thus:
$\sup_{l\in[0,\bar{l}]}|\hat{r}_{3}(l)-r_{3}(l)|\to 0.$

### v3

@@ 4.1 Uniform Convergence of Distribution
    For any given interval $\tau_{\ell},\tau_{u}]\subset(0,1)$ with $0<\tau_{\ell}<\tau_{u}<1$, the density functions $f_{Y}(\cdot\,|\,\mathbf{x})$ are uniformly bounded away from 0 and $\infty$ in the sense that
@@
    For any given interval $\tau_{\ell},\tau_{u}]\subset(0,1)$ with $0<\tau_{\ell}<\tau_{u}<1$, there exist positive definite matrices $D_{0}$ and $D_{1}(\tau)$ such that
@@
Assumption 2 and 3 are commonly adopted in quantile regression analysis to ensure robust and reliable statistical inferences. These conditions facilitate the estimation of quantile regression coefficients and provide the theoretical underpinning for the asymptotic analysis.
With the above assumptions, the following uniform convergence of $F_{\hat{Y}}(y\,|\,\mathbf{x})$ has been established by hong2023learning.
Lemma 1 (Hong et.al, 2023).
@@
    For any $\mathbf{x}\in\mathcal{X}$, $\hat{Y}(\mathbf{x})\overset{d}{\to}Y(\mathbf{x})$ as $n,m\to\infty$, where $\overset{d}{\to}$” denotes convergence in distribution.
@@
The uniform convergence of $F_{\hat{Y}}(y\,|\,\mathbf{x})$ established in Lemma 1 serves as the central tool in analyzing the asymptotic behaviors of estimators of risk measures, which will be discussed in the next subsection.
4.2 Asymptotic Analysis for Estimators of Risk Measures
Building on the uniform convergence of the conditional distribution function established in Lemma 1, we now analyze the risk measure
$r_{3}(l)\;=\;\mathbb{E}\!\left[\,g_{l}(Y(\mathbf{x}))\,\mathbb{I}\{Y(\mathbf{x})<a(l)\}\right],\quad l\in[0,\bar{l}],$
and its QRGMM-based plug-in estimator
$\hat{r}_{3}(l)\;=\;\mathbb{E}\!\left[\,g_{l}(\hat{Y}(\mathbf{x}))\,\mathbb{I}\{\hat{Y}(\mathbf{x})<a(l)\}\right],$
where we ignore Monte Carlo error from sampling $\hat{Y}(\mathbf{x})$. It is important to note that $r_{3}(l)$ is the most general form of risk measure that we consider. It includes $r_{1}(l)$ and $r_{2}(l)$ as special cases.
Two sources of difficulty arise in the analysis. (i) Expectation of a transformation. Even though $F_{\hat{Y}}(\cdot\,|\,\mathbf{x})$ converges uniformly to $F_{Y}(\cdot\,|\,\mathbf{x})$, weak convergence alone does not in general imply convergence of expectations for nonlinear transforms such as $g_{l}(y)$. (ii) Discontinuity from the indicator. The factor $\mathbb{I}\{y<a(l)\}$ renders the map $y\mapsto g_{l}(y)\mathbb{I}\{y<a(l)\}$ potentially discontinuous at $a(l)$. Consequently, standard asymptotic convergence theory for bounded continuous functions, such as the Portmanteau theorem (see, e.g., van2000, p. 6), does not directly apply.
To study the asymptotic behavior of $\hat{r}_{3}(l)$, we first impose the following mild regularity conditions on $g_{l}(\cdot)$ and $a(l)$:
@@ 4.2 Asymptotic Analysis for Estimators of Risk Measures
    $a(l)$ is monotone on $0,\bar{l}]$, satisfies $0\leq a(l)\leq a(\bar{l})<\infty$, with $a(l)\to 0$ as $l\to 0$.
@@
    $g_{l}(\cdot)$ is Lipschitz continuous with constant $K$ within $y\in[0,a(\bar{l})]$ uniformly over $l\in[0,\bar{l}]$, and uniformly bounded by $g_{\bar{l}}(0)<\infty$.
Assumption (4.a) requires $a(l)$, a threshold function, to be monotonic in $l$ and vanish as $l\to 0$, reflecting the intuition that smaller loans imply lower trigger levels. Assumption (4.b) requires that the loss function $g_{l}(\cdot)$ to be Lipschitz continuous with constant $K$ on $0,a(\bar{l})]$, which by Rademacher’s theorem (see, e.g., evans2018measure) ensures that $g_{l}(\cdot)$ is differentiable almost everywhere (a.e.) and that its derivative satisfies $\left|g_{l}^{\prime}(y)\right|\leq K$ a.e.. Moreover, $g_{l}(\cdot)$ is uniformly bounded by $M=g_{\bar{l}}(0)$, reflecting the maximum loss in the worst case when sales $Y(\mathbf{x})=0$ and loan $l=\bar{l}$. These conditions are in general intuitive and reasonable for SCF risk management. They ensure tractable analysis of $\hat{r}_{3}(l)$ and lead to the following result.
@@
Under Assumptions 14, for any given $\mathbf{x}$, as $n,m\to\infty$, we have
@@
For any $\mathbf{x}\in\mathcal{X}$ and $l\in[0,\bar{l}]$,
$r_{3}(l)$ $=\mathbb{E}\left[g_{l}(Y(\mathbf{x}))\cdot\mathbb{I}_{\{Y(\mathbf{x})<a(l)\}}\right],$
$\hat{r}_{3}(l)$ $=\mathbb{E}\left[g_{l}(\hat{Y}(\mathbf{x}))\cdot\mathbb{I}_{\{\hat{Y}(\mathbf{x})<a(l)\}}\right].$
By rewriting the expectations in integral form, we obtain
$r_{3}(l)$ $=\int_{0}^{a(l)}g_{l}(y)\,dF_{Y}(y\,|\,\mathbf{x}),$
$\hat{r}_{3}(l)$ $=\int_{0}^{a(l)}g_{l}(y)\,dF_{\hat{Y}}(y\,|\,\mathbf{x}).$
Hence,
$\hat{r}_{3}(l)-r_{3}(l)=\int_{0}^{a(l)}g_{l}(y)\,d(F_{\hat{Y}}-F_{Y})(y\,|\,\mathbf{x}).$
Since $g_{l}(\cdot)$ is Lipschitz continuous with constant $K$ on $0,a(\bar{l})]$, uniformly over $l\in[0,\bar{l}]$, it follows that $g_{l}(y)$ is differentiable a.e. on this domain, with $\left|g_{l}^{\prime}(y)\right|\leq K$ a.e.. Moreover, since $g_{l}(\cdot)$ is Lipschitz, it is absolutely continuous, so the integration by parts formula can be applied under the Lebesgue integral.
Applying integration by parts to $\hat{r}_{3}(l)-r_{3}(l)$ yields
@@
$=$ $\left.\left[g_{l}(y)\left(F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right)\right]\right|_{0}^{a(l)}$
$-\int_{0}^{a(l)}\left(F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right)g_{l}^{\prime}(y)\,dy.$
Therefore,
$\hat{r}_{3}(l)-r_{3}(l)$ $=$ $g_{l}(a(l))\left(F_{\hat{Y}}(a(l)\,|\,\mathbf{x})-F_{Y}(a(l)\,|\,\mathbf{x})\right)$
$-\ g_{l}(0)\left(F_{\hat{Y}}(0\,|\,\mathbf{x})-F_{Y}(0\,|\,\mathbf{x})\right)$
$-\int_{0}^{a(l)}\left(F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right)g_{l}^{\prime}(y)\,dy.$
Let
$\delta_{n,m}=\sup_{y\geq 0}\left|F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right|,$ (3)
and denote $M=g_{\bar{l}}(0)$. For the boundary terms, since $\left|g_{l}(a(l))\right|\leq M$, $\left|g_{l}(0)\right|\leq M$, and $\left|F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right|\leq\delta_{n,m}$, it follows that
$\left|g_{l}(a(l))\right|\cdot\left|F_{\hat{Y}}(a(l)\,|\,\mathbf{x})-F_{Y}(a(l)\,|\,\mathbf{x})\right|\leq M\delta_{n,m},$
$\left|g_{l}(0)\right|\cdot\left|F_{\hat{Y}}(0\,|\,\mathbf{x})-F_{Y}(0\,|\,\mathbf{x})\right|\leq M\delta_{n,m}.$
For the integral term, using $\left|F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right|\leq\delta_{n,m}$ and $\left|g_{l}^{\prime}(y)\right|\leq K$ a.e., we obtain
$\int_{0}^{a(l)}\left|F_{\hat{Y}}(y\,|\,\mathbf{x})-F_{Y}(y\,|\,\mathbf{x})\right|\cdot\left|g_{l}^{\prime}(y)\right|\,dy$
$\leq$ $\int_{0}^{a(l)}\delta_{n,m}K\,dy=K\delta_{n,m}a(l)\leq K\delta_{n,m}a(\bar{l}),$
as $a(l)\leq a(\bar{l})$. Combining the bounds above gives
$\left|\hat{r}_{3}(l)-r_{3}(l)\right|$ $\leq$ $M\delta_{n,m}+M\delta_{n,m}+K\delta_{n,m}a(\bar{l})$
@@
Because the bound is independent of $l$, it holds uniformly:
$\sup_{l\in[0,\bar{l}]}\left|\hat{r}_{3}(l)-r_{3}(l)\right|\leq(2M+Ka(\bar{l}))\delta_{n,m}.$ (4)
By Lemma 1(b), we have $\delta_{n,m}\to 0$ as $n,m\to\infty$. Thus,
$\sup_{l\in[0,\bar{l}]}\left|\hat{r}_{3}(l)-r_{3}(l)\right|\to 0.$